# Equilibrium and out-of-equilibrium critical dynamics of the three-dimensional Heisenberg model with random cubic anisotropy

### v1

@@ Abstract
We study the critical dynamics of the three-dimensional Heisenberg model with random cubic anisotropy in the out-of-equilibrium and equilibrium regimes. Analytical approaches based on field theory predict that the universality class of this model is that of the three-dimensional site-diluted Ising model. We have been able to estimate the dynamic critical exponent by working in the equilibrium regime and by computing the integrated autocorrelation times obtaining $z=2.50(5)$ (free scaling corrections) and $z=2.31(7)$ (by fixing the scaling corrections to that of the 3D site-diluted Ising model). In the out-of-equilibrium regime we have focused in the study of the dynamic correlation length which has allowed us to compute the dynamic critical exponent obtaining $z=2.38(2)$, which is compatible with the equilibrium ones. Finally, both estimates are also compatible with the most accurate prediction $z=2.35(2)$ from numerical simulations of the 3D site-diluted Ising model, in agreement with the predictions based on field theory.
@@ I Introduction
For the cubic anisotropy disorder, RG shows that the system behaves at criticality like the 3D site-diluted Ising model (RSIM) Dudka et al. (2005b).
@@
In Ref. Ruiz-Lorenzo et al. (2022) the (static) equilibrium critical phenomena of both isotropic and cubic quenched disorders was characterized by means of numerical simulations. For the cubic anisotropy disorder, a second order phase transition in the ferromagnetic channel raised, but with critical exponents ($\eta$ and $\nu$) not near those of the RSIM. Furthermore, in the isotropic disorder, a spin glass phase transition was found. However, the characterization of the dynamics of these models with both types of disorder is still lacking.
@@
The structure of the paper is the following. In Sec. II we describe the model and the observables used in the equilibrium and out-of-equilibrium studies. Next, in Sec. III, we report details on our estimate of the dynamic critical exponent, first in equilibrium, using the integrated autocorrelation times, and then out-of-equilibrium by computing and analyzing the dynamic correlation length. In the last section we will discuss the results and report our conclusions.
@@ II The model and Observables
In the cubic anisotropy disorder the vectors $\hat{\mathitbf n}_{\mathitbf r}$ point along the six semi-axes of the cubic lattice with the same probability $1/6$:
@@ II.1 Equilibrium
where ()delimited-⟨$\langle$($\cdots$)$\rangle$ ( ⋯ ) ⟩ is the thermal average and $\overline{(\cdots)}$ is the average over the disorder.
@@ II.2 Out of equilibrium
In this work, our focus lies on $\xi_{2,3}$. In spin glasses, $\xi_{1,2}$ was measured using a correlation function that decays as $1/r^{0.5}$ Belletti et al. (2008, 2009). However, in our case, to reduce the influence of shorter distances, we have computed higher values of $I_{k}$. We refer the reader to Refs.  Belletti et al. (2008, 2009); Astillero and Ruiz-Lorenzo (2019); Yllanes (2011); Young (2015) for a detailed description of the procedure we employed to calculate the integrals and estimate the statistical error associated with $\xi_{k,k+1}(t)$.
@@ III Numerical results
To perform the equilibrium numerical simulations we have used two different algorithms. In the first algorithm (denoted Algorithm A) the Metropolis algorithm is implemented by modifying the original spin by adding a random vector (not necessarily in the unit sphere) and normalizing the final vector to the unit sphere, finally the magnitude of the random vector is modified in order to maintain an acceptance between 40% and 60% Amit and Martín-Mayor (2005).
We have also used a second algorithm (termed as algorithm B) in which the Metropolis algorithm is carried out by proposing a random spin in the unit sphere. Both versions of the Metropolis algorithm are supposed to belong to the same dynamic universality class Sokal (1992).
@@ III.1 Equilibrium
We have analyzed the numerical data reported in Tables 1 and 2 neglecting the scaling corrections (e.g., using $\tau\propto L^{z}$). The results for the dynamic critical exponent are reported in Table 3 for the four simulated cases. Note that these $z$ estimates are near the RSIM dynamic exponent $z=2.35(2)$, but a bit smaller for $D=3$ and a bit larger for $D=4$. This results also strongly support that both algorithms belong to the same dynamic universality class.
Moreover, assuming universality, the four cases should present the same critical behavior and to show this fact, it is compulsory to take into account the corrections to scaling (the same $z$ exponent with maybe different scaling corrections depending of the strength of the different amplitudes of these subdominant terms)  Ballesteros et al. (1998); Hasenbusch et al. (2007b, a); Hasenbusch et al. (2008); Fernandez et al. (2016); Fytas and Martín-Mayor (2013); Fytas et al. (2016, 2019).
To check this scenario, we have tried a joint fit of the four cases using the same $z$ exponent but different $\omega$ for $D=3$ and $D=4$-data (denoted as $\omega_{D=3}$ and $\omega_{D=4}$) [^1] To tackle with correction-to-scaling terms with very different amplitudes., obtaining $z=2.50(5)$, $\omega_{D=3}=0.61(6)$ and $\omega_{D=4}=1.8(7)$ with $L_{\mathrm{min}}=8$, $\chi^{2}/N_{\mathrm{DOF}}=1.5$ and $N_{\mathrm{DOF}}=10$ and an associated $p$-value=0.12 (the goodness of the fit is reported via $\chi^{2}/N_{\mathrm{DOF}}$, $N_{\mathrm{DOF}}$ stands for the number of degrees of freedom of the fit, and the associated $p$-value also denoted as $Q$ Young (2015)).
We also report this fit and the numerical data for the integrated correlation time in Fig. 1.
@@
Finally, since the correction-to-scaling exponent of $D=3$ is very near the value of the RSIM, $\omega=0.82(8)$ (the sub-leading one), we can try a joint fit of the $D=3$-data (see Table 1) assuming this value obtaining $z=2.31(6)(1)$ with $\chi^{2}/N_{\mathrm{DOF}}=0.46$ and $N_{\mathrm{DOF}}=4$ and an associated $p$-value=0.76 with $L_{\mathrm{min}}=8$. The second error bar on $z$ is due to the uncertainty of the value of $\omega=0.82(8)$. Therefore a very good agreement with the RSIM dynamic exponent $z=2.35(2)$ has been obtained. The same analysis on the $D=4$-data (Table 2), however, provides a $z$ value with a bigger error ($35\%$).
@@ III.2 Out-of-equilibrium: Correlation length
Firstly, we have studied the finite size effects of the out-of-equilibrium numerical simulations. To do that, we have computed the dynamic critical correlation length, $\xi_{23}$, using $L=128$ and $L=250$ sizes. In Fig. 2 we show the difference between the correlation length ($\xi_{23}$) computed simulating both lattice sizes, and we can assess that (in the statistical errors) both estimates of $\xi_{23}$ are compatibles as far as the Monte Carlo time is less than $10^{4}$ Monte Carlo sweeps. Hence, our out-of-equilibrium analysis will be based on the $L=128$ lattice for which we have simulated a larger number of samples. We report in Table 4 the number of samples we have simulated in the $L=128$ lattice for the two values of the cubic anisotropy. [^2] In the $L=250$ simulations we have simulated 1000 samples for $D=3$ and $1400$ for $D=4$.
Figure 3 shows the behavior of the dynamic correlation length ($\xi_{23}$) as a function of time for a $L=128$ lattice and $D=3$ and $D=4$. The slope of the curves (in a double logarithmic scale) provided us with the estimate of $1/z$, neglecting scaling corrections, and we report our estimates for $z$ in Table 4. As in the equilibrium analysis, we have used the jackknife method over the samples to compute the statistical errors of the different observables, including the dynamic critical exponent Astillero and Ruiz-Lorenzo (2019).
@@
Figure 3: (color online) Behavior of $\xi_{23}$ versus time for $L=128$ and for $D=3$ and $D=4$. The straight lines are the fit, in the interval $[1000,6000]$ for the pure law behavior, see Eq. (14), the described in the text.
@@ IV Conclusions
We have obtained $z=2.50(5)$ for the equilibrium analysis doing a joint fit of the numerical behavior of the integrated correlation time for both values of $D$ and for both algorithms and taking into account scaling corrections, obtaining $\omega\sim 0.61(6)$ which is fully compatible with the sub-leading correction to scaling exponent of the RSIM (the leading scaling correction is $\omega=0.29(2)$Hasenbusch et al. (2007b, a): the computed dynamic exponent is slightly larger, 2.8 standard deviations from that of the RSIM. However, by using only the $D=3$-data and assuming the sub-leading correction-to-scaling exponent of the RSIM ($\omega=0.82(8)$) we obtain a very good agreement: $z=2.31(7)$ versus $2.35(2)$.
Notice that in the equilibrium numerical studies of the Random Anisotropy model with cubic anisotropy we have simulated vector spins (instead of Ising ones) and we have not at our disposal clusters methods (to generate the equilibrated initial conditions, which will be evolved using a local Monte Carlo algorithm, the Metropolis algorithm in our case). Finally, only assuming the theoretical prediction, we can fix the scaling corrections in the analysis of the different observables of the model.
@@
The field-theoretical approach of the renormalization group shows that the universality class of the 3D Heisenberg model with cubic anisotropy is that of the RSIM for both static and dynamic. Equilibrium numerical simulations reported static critical exponents not so compatible with those of the RSIM, $\eta=0.03(1)$ and $\nu=0.737(8)$ Ruiz-Lorenzo et al. (2022) to be confronted with those of the RSIM: $\eta=0.036(1)$ and $\nu=0.683(2)$Hasenbusch et al. (2007b). Despite this fact, the dynamic critical exponent $z$ computed in this work is compatible with that of the RSIM, $z=2.35(2)$ Hasenbusch et al. (2007a), using equilibrium and out-of-equilibrium methodologies.
Hence we have provided strong support for a dynamic critical exponent computed in equilibrium (including universality, independence of $z$ of the simulated values of $D$ and of the implementation of the Metropolis algorithm) and out-of-equilibrium near to the dynamic exponent of the RSIM, accordingly with the predictions of perturbative field theory.
However, it remains an open problem to understand the difference among the static critical exponents found in the 3D Heisenberg model with cubic anisotropy and those of the RSIM.
@@ V Acknowledgments
We thank A. Maiorano, M. Dudka and Y. Holovatch for discussions.
Financial support from Grant No. PID2020-112936GB-I00 funded by the Spanish agency MCIN/AEI/10.13039/501100011033 and from Grant No. IB20079 funded by Junta de Extremadura (Spain) and by ERDF “A way of making Europe” was acknowledged.

### v2

@@ Abstract
We study the critical dynamics of the three-dimensional Heisenberg model with random cubic anisotropy in the out-of-equilibrium and equilibrium regimes. Analytical approaches based on field theory predict that the universality class of this model is that of the three-dimensional site-diluted Ising model. We have been able to estimate the dynamic critical exponent by working in the equilibrium regime and by computing the integrated autocorrelation times obtaining $z=2.50(5)$ (without taking into account scaling corrections) and $z=2.29(11)$ (by fixing the scaling corrections to that predicted by field theory). In the out-of-equilibrium regime we have focused in the study of the dynamic correlation length which has allowed us to compute the dynamic critical exponent obtaining $z=2.38(2)$, which is compatible with the equilibrium ones. Finally, both estimates are also compatible with the most accurate prediction $z=2.35(2)$, from numerical simulations of the 3D site-diluted Ising model, in agreement with the predictions based on field theory.
@@ I Introduction
For the random cubic anisotropy disorder (RAM), RG shows that the system behaves at criticality like the 3D site-diluted Ising model (RSIM) Calabrese et al. (2004); Dudka et al. (2005b) with the scaling dimension (length) of the leading irrelevant scaling field is given by $\omega_{\mathrm{RAM}}=-\alpha_{\mathrm{RSIM}}/\nu_{\mathrm{RSIM}}$, being $\nu_{\mathrm{RSIM}}$ and $\alpha_{\mathrm{RSIM}}$ the thermal and specific heat critical exponent of the RSIM Calabrese et al. (2004).
@@
In Ref. Ruiz-Lorenzo et al. (2022) the (static) equilibrium critical phenomena of both isotropic and cubic quenched disorders was characterized by means of numerical simulations. For the cubic anisotropy disorder, a second order phase transition in the ferromagnetic channel raised, but with critical exponents ($\eta$ and $\nu$) not near those of the RSIM but without the numerical characterization of the scaling corrections. Furthermore, in the isotropic disorder, a spin glass phase transition was found. However, the characterization of the dynamics of these models with both types of disorder is still lacking.
@@
The structure of the paper is the following. In Sec. II we describe the model and the observables used in the equilibrium and out-of-equilibrium studies. Next, in Sec. III, we report details on our estimate of the dynamic critical exponent, first in equilibrium, using the integrated autocorrelation times, and then out-of-equilibrium by computing and analyzing the dynamic correlation length. In the last section we will discuss the results and report our conclusions. Finally, in an appendix, we describe some technical details on the procedure we employed to calculate some observables in the out-of-equilibrium dynamics.
@@ II The model and Observables
In the cubic anisotropy disorder the vectors $\hat{\mathitbf n}_{\mathitbf r}$ point along the six semi-axes of the cubic lattice with the same probability 1/61/61 / 6:
@@ II.1 Equilibrium
where ($\langle$($\cdots$)$\rangle$ ( ⋯ ) ⟩ is the thermal average and $\overline{(\cdots)}$ is the average over the disorder.
@@ II.2 Out of equilibrium
In this work, our focus lies on $\xi_{2,3}$. In spin glasses, $\xi_{1,2}$ was measured using a correlation function that decays as $1/r^{0.5}$ Belletti et al. (2008, 2009). However, in our case, to reduce the influence of shorter distances, we have computed higher values of $I_{k}$. We refer the reader to Refs.  Belletti et al. (2008, 2009); Astillero and Ruiz-Lorenzo (2019); Yllanes (2011); Young (2015) for a detailed description of the procedure we employed to calculate the integrals and estimate the statistical error associated with $\xi_{k,k+1}(t)$ (see Appendix A for a detailed description).
@@ III Numerical results
To perform the equilibrium numerical simulations we have used two different algorithms. In the first algorithm (denoted Algorithm A) the Metropolis algorithm is implemented by modifying the original spin by adding a random vector (chosen with uniform probability on the surface of the sphere of radius $\delta$) and normalizing the final vector to the unit sphere, the magnitude of the random vector $\delta$ is adjusted in such a way that we maintain an acceptance between 40% and 60% (see, for examble, Ref. Amit and Martín-Mayor (2005)).
We have also used a second algorithm (termed as algorithm B) in which the Metropolis algorithm is carried out by proposing a random spin in the unit sphere (chosen with uniform probability). Both versions of the Metropolis algorithm are supposed to belong to the same dynamic universality class Sokal (1992).
@@ III.1 Equilibrium
We have analyzed the numerical data reported in Tables 1 and 2 neglecting the scaling corrections (e.g., using $\tau\propto L^{z}$). The results for the dynamic critical exponent are reported in Table 3 for the four simulated cases. Note that these $z$ estimates are near the RSIM dynamic exponent $z=2.35(2)$, but a bit smaller for $D=3$ and a bit larger for $D=4$. This results also support that both algorithms belong to the same dynamic universality class.
Moreover, assuming universality, the four cases should present the same critical behavior and to show this fact, it is compulsory to take into account the corrections to scaling (the same $z$ exponent with the same scaling corrections)  Ballesteros et al. (1998); Hasenbusch et al. (2007b, a); Hasenbusch et al. (2008); Fernandez et al. (2016); Fytas and Martín-Mayor (2013); Fytas et al. (2016, 2019).
To check this scenario, we have tried a joint fit of the four cases using the same $z$ exponent and the same field-theory result for the leading irrelevant scaling field $\omega_{\mathrm{RAM}}=-\alpha_{\mathrm{RSIM}}/\nu_{\mathrm{RSIM}}=3-2/\nu_{\mathrm{RSIM}}=0.072(6)$ Calabrese et al. (2004) where we have used $\nu_{\mathrm{RSIM}}=0.683(2)$ Hasenbusch et al. (2007b): this is a very small value for the $\omega$ exponent, so we are confronting strong scaling corrections.
From this fit, we obtained $z=2.29(11)$ with $\chi^{2}/N_{\mathrm{DOF}}=18/10$ and a associated $p$-value=0.055.111The goodness of the fit is reported via $\chi^{2}/N_{\mathrm{DOF}}$, $N_{\mathrm{DOF}}$ stands for the number of degrees of freedom of the fit, and the associated $p$-value is also denoted as $Q$ Young (2015). To have an acceptable $p$-value we have used the $L\geq 8$ data for the Algorithm A and the $L\geq 12$ data for the algorithm B.
Taking into account the uncertainty in the value of $\omega$ provides an additional uncertainty of $5\times 10^{-3}$ in the value of $z$, which is neglibigle in confronting with the computed error (0.11). Hence, our final estimate is: $z=2.29(11)$.
@@ III.2 Out-of-equilibrium: Correlation length
Firstly, we have studied the finite size effects of the out-of-equilibrium numerical simulations. To do that, we have computed the dynamic critical correlation length, $\xi_{23}$, using $L=128$ and $L=250$ sizes. In Fig. 2 we show the difference between the correlation length ($\xi_{23}$) computed simulating both lattice sizes, and we can assess that (in the statistical errors) both estimates of $\xi_{23}$ are compatibles as far as the Monte Carlo time is less than $10^{4}$ Monte Carlo sweeps. Hence, our out-of-equilibrium analysis will be based on the $L=128$ lattice for which we have simulated a larger number of samples. We report in Table 4 the number of samples we have simulated in the $L=128$ lattice for the two values of the cubic anisotropy. [^2] In the $L=250$ simulations we have simulated 1000 samples for $D=3$ and 140014001400 for $D=4$.
[…]
@@
In Appendix A we report the detailed procedure we have followed in order to compute the statistical errors on the dynamic correlation length and on the $z$ exponent.
@@
Figure 3: (color online) Behavior of $\xi_{23}$ versus time for $L=128$ and for $D=3$ and $D=4$. The straight lines are the fit, in the interval [1000,6000][1000,6000][ 1000 , 6000 ] for the pure law behavior, see Eq. (14), the described in the text.
@@ IV Conclusions
We have obtained $z=2.29(11)$ for the equilibrium analysis doing a joint fit of the numerical behavior of the integrated correlation time for both values of $D$ and for both algorithms and taking into account the leading scaling correction provided by the field theoretical analysis of the model Calabrese et al. (2004) ($\omega=0.072(6)$). The comparison of the computed value with that of the RSIM is good: $z=2.29(11)$ versus 2.35(2)2.35(2)2.35 ( 2 ).
Notice we have a large statistical error (a factor 6) in confronting with that of the RSIM. It is important to remark that in the equilibrium numerical studies of the Random Anisotropy model with cubic anisotropy we have simulated vector spins (instead of Ising ones) and we have not at our disposal clusters methods (to generate the equilibrated initial conditions, which will be evolved using a local Monte Carlo algorithm, the Metropolis algorithm in our case). Finally, only assuming the theoretical prediction, we can fix the scaling correction in the analysis of the equilibrium autocorrelation times.
@@
The field-theoretical approach of the renormalization group shows that the universality class of the 3D Heisenberg model with cubic anisotropy is that of the RSIM for both static and dynamic. Equilibrium numerical simulations reported static critical exponents not so compatible with those of the RSIM, $\eta=0.03(1)$ and $\nu=0.737(8)$ Ruiz-Lorenzo et al. (2022) to be confronted with those of the RSIM: $\eta=0.036(1)$ and $\nu=0.683(2)$Hasenbusch et al. (2007b). We remark that in Ref.  Ruiz-Lorenzo et al. (2022) was not possible to characterize the leading scaling corrections, and that only the exponents for the crossing point of the two larger lattice were reported.
In summary, we have presented numerical evidences that the equilibrium dynamical behavior fit pretty well with the analytical predictions: i) our estimate of the dynamic critical exponent is compatible with that of the diluted ising model ($z=2.35(2)$ Hasenbusch et al. (2007a)) and ii) the scaling corrections could be parametrized using the results of Ref. Calabrese et al. (2004). The analysis of our out-of-equilibrium numerical simulations provides with a dynamic critical exponent compatible with that of the diluted Ising model, but, we have been unable to control the scaling corrections.
An open problem is to determine, by simulating larger lattices, whether the static exponents (for example, those obtained using the quotient method) can be described with $\omega=0.072$ Calabrese et al. (2004), and whether their extrapolation to infinite volume is compatible with the RSIM static exponents.
@@ V Acknowledgments
We thank M. Dudka, A. Maiorano, V. Martin-Mayor and Y. Holovatch for discussions.
We also thank an anonymous referee for suggestions and comments.
We acknowledge financial support from Grant No. PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033, and from Grant No. GR24022 funded by Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe.”
@@
Appendix A Computation of the dynamic correlation length and the $z$ exponent with their associated statistical error.
We outline the procedure followed to compute $\xi(t)$ and its associated dynamic critical exponent $z$Lulli et al. (2016); Belletti et al. (2008, 2009); Astillero and Ruiz-Lorenzo (2019); Yllanes (2011); Young (2015); Michael (1994). A key aspect of our approach is to avoid the use of the full covariance matrix, as it is often singular (see, e.g.,Yllanes (2011); Seibert (1994)). The method proceeds as follows:
  1. 1.
    The statistical error of $C(r,t)$ is estimated using the jackknife method over the set of initial conditions.
  2. 2.
    To compute $I_{k}$, we introduce a cutoff to ensure reliable signal-to-noise ratios in $C(r,t)$ at large distances:
    • The cutoff $\Lambda$ is determined by the condition $\sigma[C(\Lambda,t)]=4,C(\Lambda,t)$.
    • For a fixed time $t$, and over the range $r_{\mathrm{min}}<r<\Lambda$, the correlation function is fitted to the form:
      $C(r,t)=\frac{a_{1}}{r^{a_{2}}}\exp(-a_{3}r^{a_{4}})\,,$ (15)
      where $r_{\mathrm{min}}$ is the minimum distance for which a good fit to Eq.(15) is achieved, typically satisfying that the diagonal $\chi^{2}/\mathrm{d.o.f.}\sim 1$.
    • In order to compute $\xi(t)$, the integral in Eq.(12) is evaluated numerically using the measured values of $C(r,t)$ for $r<\Lambda$, and the fitted values from Eq.(15) for $\Lambda<r<L/2$.
    • The statistical error in $\xi(t)$ is again computed via the jackknife method over the initial conditions. The time interval used in the fit is chosen by requiring that the diagonal $\chi^{2}/\mathrm{d.o.f.}\sim 1$.
  3. 3.
    To compute the dynamic critical exponent, we perform the following two steps:
    • The values of $\xi(t)$ are then used to determine $z$ using the diagonal covariance matrix of the correlation length based on the jackknife statistical errors of the correlation length.
    • The associated error of $z$ is estimated from the standard deviation across jackknife blocks of the correlation length. The extraction of $z$ within each (jackkife) block employs the diagonal covariance matrix. Despite this simplification, the jackknife approach accurately captures correlations between different time points.
      It is important to note that the diagonal $\chi^{2}/\mathrm{d.o.f.}$ does not have the same statistical interpretation as the full (non-diagonal) version. As discussed in detail in Sec. B.3.3.1 of Ref. Yllanes (2011), this metric behaves as though there are fewer degrees of freedom, and therefore, conventional confidence intervals cannot be reliably computed.
      Furthermore, Ref. Lulli et al. (2016) demonstrates that the error bars obtained via this jackknife method—ignoring inter-data correlations—are essentially equivalent to those computed when such correlations are fully considered.